OpenStudy (annipuppi):

How do you do problems like this? Iron(III) oxide is formed wen irons combines with oxygen in the air how many grams of Fe2O3 are formed when 5.1g of Fe reacts completely with oxygen? 4Fe+3O2 --> 2Fe2O3

OpenStudy (anonymous):

Use ratio and proportion of real values against the value given in the question. According to your reaction 4 mol of Fe react with 3 of oxygen and in terms of mass, the ration is 4*56 as to 3*16 or 224g of Fe as to 48g of Oxygen So 5.1 g of Fe as to $ and $=5.1*48/224=1.092g of oxygen.

OpenStudy (annipuppi):

but it's asking how many grams of Fe2O3 not how many grams of oxygen...

OpenStudy (anonymous):

Oh, my mistake. Ok I got it now.

OpenStudy (anonymous):

Let us consider the same ratio process. The simplified molar ratio of Fe to Fe2O3 is 2 as to 1 or in terms of mass is 2*56 as to (56*2+3*16) or 112g as to 160g or in a simpler way of 1:n is 1g of Fe as to 1.43g of Fe2O3. So with 5.1g of Fe, you have $ of Fe2O3. Therefore $ =5.1*1.43/1=7.293g of Fe2O3.